### v3

# Predictability of Performance in Communication Networks Under Markovian Dynamics

Abstract
With the emergence of time-critical applications in modern communication networks, there is a growing demand for proactive network adaptation and quality of service (QoS) prediction. However, a fundamental question remains largely unexplored: how can we quantify and achieve more predictable communication systems in terms of performance? To address this gap, this paper introduces a theoretical framework for defining and analyzing predictability in communication systems, with a focus on the impact of observations for performance forecasting. We establish a mathematical definition of predictability based on the total variation distance between forecast and marginal performance distributions. A system is deemed unpredictable when the forecast distribution, providing the most comprehensive characterization of future states using all accessible information, is indistinguishable from the marginal distribution, which depicts the system’s behavior without any observational input. This framework is applied to multi-hop systems under Markovian conditions, with a detailed analysis of Geo/Geo/1 queuing models in both single-hop and multi-hop scenarios. We derive exact and approximate expressions for predictability in these systems, as well as upper bounds based on spectral analysis of the underlying Markov chains. Our results have implications for the design of efficient monitoring and prediction mechanisms in future communication networks aiming to provide dependable services.
Index Terms:
Predictability, Predictive QoS, Queuing System, Observable Markov Model
I Introduction
With the rapid growth of the number of mobile communication users and applications in recent decades, the capacity and data rates of wireless networks have increased significantly. This progress has been achieved incrementally through successive generations (e.g., from 3G to 5G), which have substantially boosted achievable throughput. In more recent years, the demand for time-critical applications has emerged with stringent requirements regarding the end-to-end delay of the communication system [1].
These applications, usually categorized into cyber-physical systems and human-in-the-loop applications, rely on the network’s end-to-end quality of service (QoS). In cyber-physical systems, scenarios such as vehicle to everything (V2X), industrial internet of things (IIoT), etc are considered where large packet delays may trigger system instability and jeopardize the safety of the system. In human-in-the-loop applications, such as those involving extended reality (XR) headsets, users demand a high level of responsiveness, and substantial packet delays can lead to motion sickness [2].
In response to these challenges, the community has established quality of service (QoS) criteria tailored to individual applications, which is standardized by 3GPP in 5G and continues to be refined for 6G. These criteria are defined by setting a delay target and a reliability level that limits the delay violation probability (DVP). For instance, a delay target of 10 ms with a reliability of 0.9999 has been set for industrial control applications [3, 4].
The stringent requirements of the majority of these use cases ask for proactive network adaptation approaches via predicting QoS KPIs such as end-to-end delay in advance. Proactive adaptation can offer major gains to the operation of these applications, for instance, predicting in advance a potential increase in the network delay and/or deterioration of computing resources will enable the network and/or the respective application to adapt and achieve the required quality of service [5, 6, 7, 8].
Consequently, there is a strong focus in the development of 6G on utilizing end-to-end delay probability predictors, especially in the integration with time sensitive networking (TSN) [9, 10]. Moreover, 5G Automotive Association (5GAA) has introduced the concept of predictive QoS, which refers to the mechanisms enabling mobile networks to provide notifications about predicted QoS changes to interested consumers in advance [11]. ETSI, in the context of mobile edge computing (MEC), has also introduced the notion of predictive QoS support [12]. In this context, the prediction of potential handovers leading to the estimated QoS performance is described as the key solution for mitigating any service interruption and support the application requirements.
Predictability in this context refers to the ability to anticipate the future performance of the system based on current and historical data. As the communication system interacts with the environment, the predictability of the system extends beyond just the immediate network parameters. If the environmental conditions—such as interference patterns, user mobility, or traffic loads—exhibit non-random behavior and can be forecasted with reasonable accuracy, this adds a new dimension to the system’s overall predictability. Therefore, it is essential to understand how a communication system can be considered predictable, and to identify the factors that enhance or undermine this predictability.
Inspired by the above questions, in this paper, we establish a mathematical definition of predictability based on information theory concepts and a probabilistic forecasting setup. We introduce this definition and examine its consequences within a multihop communication system model where the conditions are governed by a Markov chain. We study both exact and approximate derivations of predictability within these abstract systems, such as a queueing system, for end-to-end delay prediction. Queuing systems can effectively model large-scale effects in communication networks. Our primary contributions are summarized as follows:
  • The paper introduces a mathematical definition of predictability based on probabilistic forecasting problem. A system is considered unpredictable if the probability distribution of performance conditioned on the observations (forecast distribution) is not different from the marginal distribution of performance. This is done using the total variation distance between the forecast and marginal distributions to quantify the degree of predictability.
  • We applied this framework in an abstract model with multihop systems where their performance is modulated by Markov chains and derive both exact and approximate measures of predictability. Our analysis explored the impact of system observability, state aggregation, and delayed observations on predictability.
  • We utilized spectral analysis techniques to understand how the dynamics of Markovian conditions and the randomness of their transitions influence predictability. We derive upper bounds on predictability in Markov-modulated systems via spectral analysis and the spectral gap of the Markov chain.
  • The predictability derivations are applied to an abstract queuing system, focusing on a Geo/Geo/1 queue in both single and multi-hop scenarios. In this analysis of end-to-end delay, queue length observations are considered the primary predictive factor. The study includes exact and approximate derivations of predictability.
  • Finally, by numerical evaluations, we showcased the utility of our predictability analysis in the queuing setup. The evaluations investigate the trade-offs between predictability and observability in single and multi-hop queueing systems. We demonstrated the accuracy of our proposed predictability measures and approximations by comparing them with exact calculations.
The rest of the paper is organized as follows: Section 2 presents an overview of the existing work in the literature on predictability. Section 3 presents the system model and predictability definitions, Section 4 presents the predictability analysis in Markov-modulated systems in general, while Section 5 contains the numerical evaluations and assessment of results. Finally, Section 5 concludes this work.
II Related Works
Researchers have developed various measures to assess the predictability of complex dynamic systems, all of which focus on predicting specific single or multi-dimensional critical metrics. These works can be categorized based on the following criteria: the comprehensiveness of the predictability definition, whether they analyze only the target metric or incorporate side information, and whether predictability is derived empirically or through a model-driven approach.
At the most basic level, there are works that define predictability in a setup where point prediction of a univariate metric is desired. Autocorrelation which evaluates how well a time series correlates with its own time-lagged versions, is a widely used indicator of predictability in finance [13]. Permutation entropy (PE) is the metric employed in fields like ecology and physics, to gauge predictability [14, 15, 16]. PE measures time series complexity by quantifying the diversity and frequency of ordinal patterns formed by subsequences of the series. A higher permutation entropy indicates more diverse motifs, leading to lower predictability. This indicator is mostly empirically evaluated and has been applied to various domains, including predicting infectious disease outbreaks [17], and ecological systems [16]. Abeliuk et al. explore the relationship between predictability and sampling in partially observed systems. They evaluate multiple predictability measures such as PE and autocorrelation in outbreaks of infectious diseases, digital forum interactions, and software development projects [18].
Besides autocorrelation and PE, another common approach in defining predictability is to find the lower bound on prediction error [19]. In some works, the empirically determined entropy of the target metric is incorporated in Fano’s inequality to derive the bound on error. Using this approach Li et al. investigate the limits of predictability for urban vehicular location and staying time [20]. More information-theoretic measures of predictability have inspired numerous studies. For instance, the studies on the predictability of human mobility and conversation use random entropy, uncorrelated entropy, and conditional entropy to measure predictability [21, 22].
More advanced studies in this category incorporate side information or covariates into the predictability model. As a result, the selected predictability measure is typically relative entropy, conditional entropy, or mutual information between the side information and the target metric. Bialek et al. define predictive information as the mutual information between past and future states of a time series as a criterion for predictability [23]. Haven et al. used a Gaussian ensemble prediction to assess relative entropy for predictability analysis [24]. More recently, Li et al. applied conditional entropy to traffic forecasting, highlighting the limitations of traffic predictability [25]. Also, Fang and Lee explore the predictability in the context of regression problems in machine learning. They analyze predictability in this context using the conditional entropy between the features and the label and introduce lower and upper bounds for it [26]. Due to the incorporation of side information, these predictability analysis are considered more comprehensive in our perspective.
In atmospheric sciences literature, where the probabilistic forecasting of the weather systems is of interest, the information-theoretic definition of predictability with side information is gaining more attention [27, 28]. In a series of papers, DelSole et al. define predictability as the difference between forecast and climatological distributions (marginal distribution), with the latter serving as the baseline for probabilistic information about the metric where no side information is incorporated. Various information theory measures were proposed and analyzed to quantify predictability, in systems with linear stochastic models and Gaussian processes. [29].
Among the works that are related to networking and communication systems, Ding et al. examine the predictability of radio spectrum state dynamics using Fano’s inequality. Their findings show high predictability in real-world spectrum measurements, which has significant implications for cognitive radio networks and 5G spectrum sharing [30]. In another paper, Sang et al. apply predictability analysis to an abstract network model with a queue. They investigate the predictability of network traffic by analyzing how far into the future traffic rates can be predicted with a given error constraint on the prediction problem [31].
Despite the growing focus on quality of service prediction in communication systems, comprehensive predictability analyses are both necessary and, to the best of our knowledge, largely nonexistent. This work addresses this gap by adopting a thorough predictability model and demonstrating its implications for communication system performance. Drawing inspiration from the predictability framework used in atmospheric science, which is an information-theoretic measure well-suited for probabilistic forecasting, our definition compares the conditional distribution of performance given the side information or observations with its marginal distribution. Such definition offers great capabilities not only in applying it to various dynamical systems, but also we show how it can model imperfect observability. Moreover, since Markov chains are widely used in queue modeling and various fields—including stochastic processes and machine learning—due to their effectiveness in modeling dynamic systems with memoryless state transitions, we employ the Markovian assumption for the system’s dynamics. In the following section we step into system model and predictability definition in more detail.
III System Model and Problem Formulation
[…]
Figure 1: Multi-hop communication system model with observable measures being conditions and performance.
Forecasting performance in a communication network becomes more manageable and precise by breaking down the system into independent modules and modeling each individually. This also allows for more efficient proactive adjustments, as we can more accurately pinpoint the sources of performance variations or declines [32]. These components can correspond to tandem data transmission links or, within a single link, to multiple processes that sequentially handle data bits to ensure reliable and efficient communication. Furthermore, the condition or state of each subsystem can be monitored and, according to domain knowledge, correlated with performance.
We consider a communication system that consists of $M$ independent subsystems or hops as shown in Figure 1. In time slot $n$ (time is slotted), subsystem $m$ exhibits a performance measure denoted by $Z^{(m)}_{n}$ which is a random variable over the discrete space $\mathcal{Z}$, for example, packet delay. $\Pr(Z^{(m)}_{n})$ denotes the marginal distribution (or prior distribution) and is considered stationary. It constitutes a baseline for the information content about performance, disregarding all observations. This distribution can be obtained by random sampling of the performance. In case of end-to-end delay (without packet drops) total performance is obtained via $Z_{n}=\sum_{m=1}^{M}Z_{n}^{(m)}$.
Moreover, each subsystem has an internal state or condition measure denoted by $X^{(m)}_{n}$ that influences performance, for example, the congestion level in the subsystem. Conditions are defined over a discrete finite state space $\mathcal{X}^{(m)}$ of size $K^{(m)}$. Then, the collection of all the state variables belonging to each subsystem is denoted by $X_{n}=\{X^{(1)}_{n},...,X^{(m)}_{n},...,X^{(M)}_{n}\}$ with size $K$. Observations of state measures across the system collected up to time $n$ are denoted by $O_{0:n}$. Depending on the scenario, observations can include all, parts, or an imperfect version of the conditions.
We consider three different observability defects in our model: first, delayed observations, which is particularly crucial when a network of remote nodes is involved in the packet route. In this scenario, monitoring information would arrive with a delay to the prediction and adaptation point, i.e., $O_{n}=X_{n-d}$. In the second case, we may choose to ignore certain subsystems and not observe their state, i.e., $O_{n}\subset X_{n}$ due to the overhead that the monitoring task introduces. The third case, which also concerns overhead, mandates aggregated states. In this scenario, we assume states of $\mathcal{X}$ are aggregated to $\mathcal{A}$, i.e., there exists a surjective map $\phi:\mathcal{X}\rightarrow\mathcal{A}$ with $|\mathcal{A}|\leq|\mathcal{X}|$ (or $\bar{K}\leq K$ as $|\mathcal{A}|=\bar{K}$). Then, the observations are coduncted on the aggregated states $O_{n}=\phi(X_{n})$.
The distribution of performance right after observations become available denoted by $\Pr(Z_{n}\mid O_{0:n})$ is the posterior distribution [33]. The marginal distribution is related to the posterior distribution by the chain rule as it is obtained by integrating out all observations:
$\Pr(Z_{n})=\sum_{o_{0:n}}\Pr(Z_{n}\mid O_{0:n}=o_{0:n})\Pr(O_{0:n}=o_{0:n}).$ (1)
The primary objective of any performance predictor in this setting is to capture the relationship between the observations and future performance. The probability distribution of the performance metric $L$ timeslots into the future, based on the observations collected up to the present moment, is the forecast distribution:
$\Pr(Z_{n+L}\mid O_{0:n}=o_{0:n}).$ (2)
We examine performance in communication systems that their conditions adhere to the Markovian property, which is applicable in numerous contexts and mathematically manageable. We define two sets of conditional independence assumptions on the system: first, $Z_{n}$ is independent of all other condition states except the current one $X_{n}$ and second, $X_{n}$ is independent of $X_{1},...,X_{n-2}$ given $X_{n-1}$ (the first-order Markov property), i.e.,
$\Pr(Z_{n}\mid Z_{0:n-1},X_{0:n})=\Pr(Z_{n}\mid X_{n}).$ (3)
A third assumption for the remainder of this work is that the simplified posterior distribution mentioned above is stationary, meaning that
$\forall L,\Pr(Z_{n+L}\mid X_{n+L})=\Pr(Z_{n}\mid X_{n}).$ (4)
For ease of notation, we represent the probability mass function (pmf) of this distribution as $\Pr(Z_{n}=z\mid X_{n}=x)=r_{x}(z)$, where $z\in\mathcal{Z}$ and $x\in\mathcal{X}$.
As shown in Figure 2, this system forms an observable Markov model (OMM) where the sequence $\{X_{0},X_{1},...,X_{n}\}$ is a discrete-time, time-homogeneous Markov chain with finite state space $\mathcal{X}$. The transition probability from state $x$ to state $y$ is denoted by $P(x,y)=\Pr(X_{n+1}=y\mid X_{n}=x)$ which forms the transition probability matrix $P$. $L$ step state transition probability is dervied via $P^{L}(x,y)$. Moreover, we assume that $X$ is an irreducible and aperiodic Markov chain, therefore it has an stationary distribution $\pi$ that state probabilities converge to for large lead times. Then, for all $x,y$ we have $P^{L}(x,y)\rightarrow\pi(y)$ as $L\rightarrow\infty$. Also, we assume the chain is reversible, which is an assumption satisfied by all random walks on undirected graphs, as well as many other Markov chains such as queues:
$\forall(x,y)\in\mathcal{X}^{2},\text{ }\pi(x)P(x,y)=\pi(y)P(y,x).$ (5)
Reversibility allows applying spectral analysis on the speed of convergence that we use later on. In reversible chains, the spectral theorem guarantees that $P$ admits $K=|\mathcal{X}|$ real eigenvalues, which can be ordered as
$1=\lambda_{1}\geq\lambda_{2}\geq...\geq\lambda_{K}>-1,$ (6)
and there is an orthonormal basis of real-valued eigenfunctions $(f_{j})_{j\leq K}$ corresponding to these eigenvalues with respect to the inner product $\langle\cdot,\cdot\rangle_{\pi}$ [34]. Closed-form solutions for eigenvalues and eigenfunctions exist for numerous basic and well-organized Markov chain models, including random walks on circular structures, multi-dimensional cubes, and various graph types.
In this paper, we will examine more specialized Markov chains, particularly the Geo/Geo/1/K queue, which represents the discrete-time counterpart of the M/M/1/K queue. In this case, the packets arrive according to a Bernoulli process and the probability of arriving during a time slot is $\lambda=1-\bar{\lambda}$. Service times are geometrically distributed, and the probability of completion of a service during a slot is $\mu=1-\bar{\mu}$. The utilization factor is denoted by $\rho=\lambda/\mu$. The performance metric $Z_{n}$ will be the sojourn time of the queue, and the queue size will be the measurable state of the system $X_{n}$. As the number of states in our analysis should be finite, we consider a maximum length of $K$ for the queue.
To formulate predictability in general, we rely on the forecast and the marginal distributions. An event is considered unpredictable not because its distribution is entirely unknown, but because the forecast distribution doesn’t significantly differ from the marginal distribution. Based on that, we define predictability in the following definition.
Definition 1
A system is unpredictable if the forecast and marginal distributions do not differ if
$\Pr(Z_{n+L}\mid O_{0:n}=o_{0:n})=\Pr(Z_{n+L}).$ (7)
According to this definition, the loss of predictability occurs when the future state $Z_{n+L}$ is statistically independent of the observations. Equivalently, we define the degree of predictability as total variation distance between the forecast and marginal distributions
$D_{n}(L)=\lVert\Pr(Z_{n+L}\mid O_{0:n}=o_{0:n})-\Pr(Z_{n+L}))\rVert_{\text{TV}},$ (8)
where the Total Variation distance (TV) is a statistical metric distance defined by
$\text{TV}(p,q)$ $:=\sup_{B\subset\mathcal{Z}}|p(B)-q(B)|$ (9)
$=\frac{1}{2}\sum_{z\in\mathcal{Z}}|p(z)-q(z)|.$ (10)
for probability mass functions $p$ and $q$.
Total variation distance is a metric that ranges between zero and one and benefits from numerous properties for bounding and approximation. Intuitively, the system is unpredictable when the TV distance between the forecast and the marginal is very small, so the performance and observations are almost independent. On the contrary, the system is predictable when TV is close to 1 since the distributions are very different and that means the observations make a difference in our prediction of the performance. This framework implies that predictability is a combined property of the dynamical system and the observations. The observations might be irrelevant, leading to zero predictability, or the system conditions might change in a completely random manner, rendering the incorporation of the observations ineffective.
It is important to acknowledge that the predictability framework introduced in this paper may not be universally applicable to all systems. Specifically, it might seem counterintuitive when applied to deterministic or nearly deterministic processes that do not need side information. For example, a process that always takes the value 1 would be considered unpredictable according to this definition because the forecast and marginal distributions are identical, yielding a total variation distance of zero. However, in such a scenario, alternative measures like entropy could be more appropriate, as they would correctly identify the process as fully predictable due to its zero entropy. Thus, the proposed definition is particularly well-suited for systems where the primary concern is the utility of observations in reducing uncertainty. This is especially relevant for networked systems where the decision to incorporate additional signals in performance prediction must be weighed against the cost of processing and routing them.
Definition 2
Epsilon-predictable horizon of a system is defined by the maximum lead time that predictability remains greater than $\epsilon$:
$H_{\epsilon}=\max\{L,\forall n:D_{n}(L)\geq\epsilon\}.$ (11)
This metric guarantees that if we have observations up to time $n$, then we will have informative forecasts up to time $n+L$, since the forecast distribution will always have minimum distance $\epsilon$ with the marginal distribution. Essentially, it reflects how quickly the observations lose their importance, as the system evolves quickly to its stationary state. Therefore, epsilon-predictable horizon can be essential in analyzing predictability of complex dynamic systems.
$X_{n-1}$ $X_{n}$ $X_{n+1}$ $Z_{n-1}$ $Z_{n}$ $Z_{n+1}$$\dots$
Figure 2: Dependencies in the performance model
In addressing the problems considered in this study, we begin by assessing the predictability of system models under conditions of imperfect observations. Specifically, we aim to determine the predictability of the systems with the following forecast distributions:
  • with delayed observation
    $\Pr(Z_{n+L}\mid O_{n}=X_{n-d}),$ (12)
  • with the observation of aggregated states
    $\Pr(Z_{n+L}\mid O_{n}=\phi(X_{n})),$ (13)
  • and with partial observation in multihop cases
    $\Pr(Z_{n+L}\mid O_{n}\subset\{X_{n}^{(1)},\cdots,X_{n}^{(M)}\}).$ (14)
Next, we will determine upper bounds on predictability based on $L$ for both single-hop and multi-hop system models, which can then be converted to an upper limit on $H_{\epsilon}$. Finally, we will derive exact and approximation solutions for the predictability of sojourn time in basic queuing systems, specifically Geo/Geo/1/K.
We will evaluate the results in both single-hop and multihop scenarios under various configurations with imperfect observations and state aggregation. Building on the system model and predictability definitions established in this section, we now delve into a detailed predictability analysis for Markov-modulated systems in the following section.
IV Predictability Analysis
To begin, we consider perfect observation of conditions, as $O_{n}=X_{n}$ for all $n$, until otherwise specified. In the following lemma, we introduce two immediate conclusions which apply to any Markov-modulated probabilistic performance measure.
Lemma 1
In a system where the performance $Z_{n}$ is related to Markov chain conditions $X_{n}$ via a known stationary posterior distribution $\Pr(Z_{n}\mid X_{n})$, forecast distribution (distribution of performance $L$ time slots ahead, given the observed state $x$) can be derived as
$\begin{split}\Pr(Z_{n+L}\mid X_{n}=x)&=\sum_{y\in\mathcal{X}}P^{L}(x,y)\Pr(Z_{n}\mid X_{n}=y)\end{split}$ (15)
Marginal distribution of performance can be obtained via
$\begin{split}\Pr(Z_{n+L})&=\sum_{y\in\mathcal{X}}\pi(y)\Pr(Z_{n}\mid X_{n}=y).\end{split}$ (16)
Proof: we begin by expanding forecast definition using law of total probability as
$\Pr(Z_{n+L}=z\mid X_{n}=x)=\\ \sum_{y\in\mathcal{X}}\Pr(Z_{n+L}=z\mid X_{n+L}=y,X_{n}=x)\\ \Pr(X_{n+L}=y\mid X_{n}=x).$ (17)
Then, according to first-order Markov property:
$\Pr(Z_{n+L}=z\mid X_{n}=x)=\\ \sum_{y\in\mathcal{X}}\Pr(Z_{n+L}=z\mid X_{n+L}=y)\\ \Pr(X_{n+L}=y\mid X_{n}=x).$ (18)
The definition of Markov chain’s transition matrix lets us simplify more:
$\Pr(Z_{n+L}=z\mid X_{n}=x)=\\ \sum_{y\in\mathcal{X}}\Pr(Z_{n+L}=z\mid X_{n+L}=y)P^{L}(x,y)=\\ \sum_{y\in\mathcal{X}}r_{y}(z)P^{L}(x,y).$ (19)
For the marginal distribution, again we expand it using law of total probability as
$\Pr(Z_{n+L}=z)=\sum_{y\in\mathcal{X}}\Pr(Z_{n+L}=z\mid X_{n+L}=y)\\ \Pr(X_{n+L}=y).$ (20)
Moreover, we use the definition of Markov chain’s stationary state probabilities to finish the proof
$\Pr(Z_{n+L}=z)=\\ \sum_{y\in\mathcal{X}}\Pr(Z_{n+L}=z\mid X_{n+L}=y)\pi(y)=\sum_{y\in\mathcal{X}}r_{y}(z)\pi(y).$ (21)
This lemma indicates that in such a system, forecast distribution becomes a mixture model with kernels defined by the posterior distribution and weights defined by Markov chain’s transient state probabilities. As for the marginal distribution, regardless of the observations, it also becomes a mixture model with the same kernels but weights are defined by the Markov chain’s stationary state probabilities
Theorem 1
Predictability of a Markov-modulated process $Z_{n}$ with Markov chain probabilities {$P$, $\pi$} and posterior distributions $r_{y}(z)$, based on the perfect observation of the state $x$ is characterized as follows:
$D_{n}(L)$ $=\sup_{A\subset\mathcal{Z}}\big{\lvert}\sum_{z\in A}\sum_{y\in\mathcal{X}}(P^{L}(x,y)-\pi(y))r_{y}(z)\big{\rvert},$ (22)
$=\frac{1}{2}\sum_{z\in\mathcal{Z}}\lvert\sum_{y\in\mathcal{X}}(P^{L}(x,y)-\pi(y))r_{y}(z)\rvert.$ (23)
Proof: Substituting the forecast and marginal distributions in lemma 1 into the predictability definition 1 results the two forms for predictability derivation.
This theorem provides an inaugural methodology for deriving the predictability of a general Markovian system. It requires characterization of the conditions as a Markov model to determine $\{P,\pi\}$ and the derivation of the conditional distribution of the target metric across all conditions $r_{y}(z)$ $\forall y$ or the posterior distribution. In data-driven scenarios, these tasks correspond to Markov model parameter estimation and conditional density estimation, respectively.
In the following subsections, we address more specific scenarios with imperfect observations, spectral analysis, multihop, and queueing systems.
IV-A Imperfect Observations Impact
We analyze the impact of imperfect observations on predictability in two parts:
Delayed Observations
This scenario is particularly critical given a network of remote nodes involved. Observations must be transmitted to the forecast location, introducing potential delays where $O_{n}=X_{n-d}$.
Lemma 2
The predictability of a Markov-modulated process for the lead time $L$ given observations delayed by $d$ timeslots equals the predictability of the system with perfect observations but for the lead time $L+d$.
Proof: Suppose an observation arrives at the forecast point with a delay of $d$ timeslots. The forecast distribution for this scenario becomes $\text{Pr}(Z_{n+L}\mid X_{n-d})$. Given that the Markov chain and posterior distribution are time invariant, we can substitute the current time index $n$ with $n+d$ as follows
$\text{Pr}(Z_{n+L}\mid X_{n-d})=\text{Pr}(Z_{n+L+d}\mid X_{n}).$ (24)
Therefore, $L+d$ can be seen as the new forecast lead time while other variables in Theorem 1 remain the same.
When the observation delay increases, it eventually exceeds the epsilon-predictable horizon, and the forecast converges to the marginal distribution, rendering the observation ineffective. Therefore, the delay in the observations must be smaller than the epsilon-predictable horizon of the system to have useful predictions.
Aggregated States Observation
In this scenario, the aggregation results in a new Markov chain with probabilities $\{\bar{P},\bar{\pi}\}$ and a new posterior distribution $\bar{r}_{a}(z)$ for any $a\in\mathcal{A}$. Since Lemma 1 and Theorem 1 would hold true for this new aggregated setup, we can derive predictability accordingly. In the following lemma, we demonstrate how to construct $\bar{P}$, $\bar{\pi}$, and $\bar{r}_{a}(z)$.
Lemma 3
Suppose the states of a Markov chain $\mathcal{X}$ are aggregated via the surjective map $\phi$ resulting a new Markov chain with state space $\mathcal{A}$ and probabilities $\{\bar{P},\bar{\pi}\}$. The aggregate transition matrix denoted by $\bar{P}$ is derived by
$\bar{P}(a,b)=\frac{\sum_{x\in\phi^{-1}(a)}\sum_{y\in\phi^{-1}(b)}P(x,y)\pi(x)}{\pi(\phi^{-1}(a))}.$ (25)
The posterior distribution is obtained via
$\bar{r}_{a}(z)=\frac{\sum_{y\in\phi^{-1}(a)}\pi(y)r_{y}(z)}{\pi(\phi^{-1}(a))}.$ (26)
The aggregate stationary distribution is evident to be
$\bar{\pi}(b)=\pi(\phi^{-1}(b)).$ (27)
Proof: We begin with expanding the transition probability matrix using the law of total probability:
$\bar{P}(a,b)=\Pr(\phi(X_{n+1})=b\mid\phi(X_{n})=a)=\\ \frac{\sum_{x\in\phi^{-1}(a)}\Pr(\phi(X_{n+1})=b\mid X_{n}=x)\Pr(X_{n}=x)}{\Pr(\phi(X_{n})=a)}.$ (28)
Then, using additivity we can substitute the numerator as
$\sum_{x\in\phi^{-1}(a)}\Pr(\phi(X_{n+1})=b\mid X_{n}=x)\Pr(X_{n}=x)=\\ \sum_{x\in\phi^{-1}(a)}\sum_{y\in\phi^{-1}(b)}\Pr(X_{n+1}=y\mid X_{n}=x)\Pr(X_{n}=x),$ (29)
and use Markov chain notations to achieve Equation 25. For the posterior distribution we apply simillar approach and obtain
$\bar{r}_{a}(z)=\Pr(Z_{n}=z\mid\phi(X_{n})=a)=\\ \frac{\sum_{y\in\phi^{-1}(a)}\Pr(Z_{n}=z\mid X_{n}=y)\Pr(X_{n}=y)}{\Pr(\phi(X_{n})=a)},$ (30)
where by substituting notations we get Equation 26.
In the next section we opt for a more intuitive and robust understanding of predictability by analyzing the inherent properties of the Markov chain with spectral analysis, rather than solely relying on the estimation/assumption of state transition probabilities.
IV-B Markov Chain Spectral Properties Impact
In the following lemma, we summarize the literature on spectral analysis which is applicable on the total variation distance to equilibrium in reversible Markov chains.
Lemma 4 ([34, 35])
In a reversible Markov chain, for any starting state $x$, the total variation distance to stationary state probabilities is bounded as
$\left\|P^{L}(x,\cdot)-\pi\right\|_{\text{TV}}\leq\frac{1}{2}\left(\sum_{j=2}^{K}\lambda^{2L}_{j}f^{2}_{j}(x)\right)^{1/2}.$ (31)
Suppose $\lambda_{*}=\text{max}\{|\lambda|,\lambda\neq 1\}$, then $\xi=1-\lambda_{*}$ denotes the spectral gap which can simplify the bound further:
$\left\|P^{L}(x,\cdot)-\pi\right\|_{\text{TV}}\leq\frac{1}{2}\lambda_{*}^{L}\left(\frac{1}{\pi(x)}-1\right)^{1/2}.$ (32)
Spectral gap is known to provide a crucial measure on the rate at which the chain converges to its stationary distribution. This measure quantifies the chain’s mixing properties: a larger spectral gap implies faster convergence, meaning that the distribution of the chain’s states will quickly approach the stationary distribution.
By employing the spectral analysis principles, we derive an upper bound on the predictability of general Markovian systems which is introduced in the following theorem.
Theorem 2 (Spectral-Based Predictability Bound)
Predictability of a Markov-modulated process $Z_{n}$ with Markov chain probabilities {$P$, $\pi$} and stationary posterior distribution $r_{y}(z)$, based on the perfect observation of the state $x$ is upper-bounded as follows:
$D_{n}(L)\leq\frac{1}{2}\left(\sum_{j=2}^{K}\lambda^{2L}_{j}f^{2}_{j}(x)\right)^{1/2}\sqrt{2}\left(R-1\right)^{1/2},$ (33)
Where
$R=\sum_{z}\frac{\sum_{y}\pi(y)r^{2}_{y}(z)}{\sum_{y}{\pi(y)r_{y}(z)}},\text{ and }1\leq R\leq K.$ (34)
The inequality can be more simplified using the second largest eigenvalue as
$D_{n}(L)\leq\frac{1}{2}\lambda^{L}_{*}\left(\frac{1}{\pi(x)}-1\right)^{1/2}\sqrt{2}\left(R-1\right)^{1/2}.$ (35)
Proof: see appendix A.
This theorem suggests that the predictability upper bound decays geometrically, with the same rate that Markov chain converges to stationarity. Upper bounds in Equations 35 and 32 are different in only one term which is sensitive to the closeness of posterior distributions $r_{y}(z)$. The minimum value of $R$, i.e. $1$, occurs when all conditional distributions are similar and makes the predictability zero as was expected. On the other hand, we get the maximum, i.e. $K=|\mathcal{X}|$, when conditional distributions have minimal overlap.
Remark 1
For the cases that $\frac{3}{2}<R\leq K$, a tighter upper bound on the predictability is the Markov chain’s total variation distance to equilibrium.
Proof: using triangle inequality and $0\leq r_{y}(z)\leq 1\quad\forall y\in\mathcal{X},\forall z\in\mathcal{Z}$ on Theorem 1 we have
$D_{n}(L)=\frac{1}{2}\sum_{z\in\mathcal{Z}}\bigg{|}\sum_{y\in\mathcal{X}}\left(P^{L}(x,y)-\pi(y)\right)r_{y}(z)\bigg{|}\\ \leq\frac{1}{2}\sum_{z\in\mathcal{Z}}\sum_{y\in\mathcal{X}}\left(\left|P^{L}(x,y)-\pi(y)\right|r_{y}(z)\right)\\ =\frac{1}{2}\sum_{y\in\mathcal{X}}\left|P^{L}(x,y)-\pi(y)\right|\sum_{z\in\mathcal{Z}}r_{y}(z)\\ =\frac{1}{2}\sum_{y\in\mathcal{X}}\left|P^{L}(x,y)-\pi(y)\right|=\frac{1}{2}\|P^{L}(x,\cdot)-\pi\|_{\text{TV}}.$ (36)
Theorem 2 introduces a novel approach to analyzing observable Markov models that modulate a secondary process through a known, stationary posterior distribution. Unlike traditional Markov chain mixing time analyses that focus on the total variation distance between stationary and transient state probabilities, our work provides an upper bound on the total variation distance between the secondary processes modulated by these different state probabilities. This advancement offers a new perspective on the convergence behavior of complex, multi-layered stochastic systems.
The subsequent section explores techniques for analyzing predictability in multi-hop systems.
IV-C Predictability in Multi-Hop Systems
End-to-end performance predictability analysis in a multi-hop system as shown in Figure 1 requires an assumption on the performance measure and how it is obtained in total. In scenarios involving multiple hops, we consider $Z_{n}$ to represent the end-to-end delay, which is determined as $Z_{n}=\sum_{m=1}^{M}Z_{n}^{(m)}$. Due to the summation, total delay distribution will be the convolution of each hop’s delay distribution both for forecasts and marginals as
$\Pr(Z_{n+L}\mid X_{n}=x)=\\ \Pr(Z^{(1)}_{n+L}\mid X^{(1)}_{n}=x^{(1)})\ast\Pr(Z^{(2)}_{n+L}\mid X^{(2)}_{n}=x^{(2)})\ast\\ \dots\ast\Pr(Z^{(M)}_{n+L}\mid X^{(M)}_{n}=x^{(M)}),$ (37)
$\Pr(Z_{n+L})=\\ \Pr(Z^{(1)}_{n+L})\ast\Pr(Z^{(2)}_{n+L})\ast\dots\ast\Pr(Z^{(M)}_{n+L}),$ (38)
where for each hop, marginal and forecast distributions can be derived via Lemma 1.
Obtaining predictability using the above method becomes less tractable as the dimensions of the system grow. In this context, identifying approximation or bounding techniques applicable to the problem is helpful. Below we propose a key lemma regarding the predictability in tandem systems where the end-to-end delay is subject to forecast.
Lemma 5 (Subadditivity of Predictability)
The predictability of a tandem multi-hop system, assuming independence of the hops, is upper-bounded via the sum of predictability of each hop as
$D_{n}(L)\leq\sum_{m=1}^{M}D^{(m)}_{n}(L).$ (39)
Proof: For independent random variables $F^{1}$, $F^{2}$, $M^{1}$, and $M^{2}$ and their distributions, the total variation distance is subadditive:
$\lVert(F^{1}+F^{2})-(M^{1}+M^{2})\rVert_{\text{TV}}\leq\lVert F^{1}-M^{1}\rVert_{\text{TV}}+\lVert F^{2}-M^{2}\rVert_{\text{TV}}.$ (40)
The proof completes by supposing $F^{i}$ and $M^{i}$ denote forecast and marginal distributions of hop $i$.
This lemma proves beneficial especially when examining the predictability of a multihop system with incomplete observability i.e. $O_{n}\subset X_{n}$. In deriving the bound by adding the predictability across all hops, if the state in one or more hops is unobserved, their predictability drops to zero and does not affect the summation in Equation 39. We will investigate this in the numerics section in more detail.
Next, we opt to apply the predictability analysis to Geo/Geo/1 queue, which is the discrete time analogue of the M/M/1 queue.
IV-D Geo/Geo/1/K Queue Predictability
In this subsection, we utilize the previously mentioned predictability concepts and methods to the Geo/Geo/1/K queue model. The target metric $Z_{n}$ for this forecast system is the sojourn time of the queue. The transition matrix in this queue is given by
$\Pr(X_{n}=j\mid X_{n-1}=i)=\\ \begin{cases}\lambda\bar{\mu}&\text{ if }j=i+1,0\leq i\leq K-1\\ \mu\bar{\lambda}&\text{ if }j=i-1,1\leq i\leq K\\ 1-\lambda\bar{\mu}&\text{ if }j=i=0\\ \lambda\mu+\bar{\lambda}\bar{\mu}&\text{ if }j=i,1\leq i\leq K-1\\ 1-\mu\bar{\lambda}&\text{ if }j=i=K\\ 0,&\text{ otherwise }\end{cases}$ (41)
and the corresponding Markov chain is depicted in Figure 3.
00 $1$dots $K$ $1-\lambda\bar{\mu}$$\lambda\bar{\mu}$$\lambda\mu+\bar{\lambda}\bar{\mu}$$\lambda\bar{\mu}$$\mu\bar{\lambda}$$\lambda\mu+\bar{\lambda}\bar{\mu}$$\lambda\bar{\mu}$$\mu\bar{\lambda}$$1-\mu\bar{\lambda}$$\mu\bar{\lambda}$
Figure 3: Geo/Geo/1/K Markov chain states and transition probabilities
In the following theorem we describe an exact derivation of predictability for this queuing model based on the observation of queue length and Theorem 1.
Theorem 3
The predictability of delay in a Geo/Geo/1/K queue given the observation of the system size denoted by $x$ is described as follows:
$D_{n}(L)=\\ \frac{1}{2}\sum_{z}\biggl{|}\sum_{y=1}^{K}\frac{2}{K+1}\beta^{\frac{y-x}{2}}\sum_{k=1}^{K}\frac{\alpha_{k}^{L}}{\left[1-2\sqrt{\beta}\cos\left(\frac{k\pi}{K+1}\right)+\beta\right]}\\ \left[\sin\left(\frac{xk\pi}{K+1}\right)-\sqrt{\beta}\sin\left(\frac{(x+1)k\pi}{K+1}\right)\right]\\ \left[\sin\left(\frac{yk\pi}{K+1}\right)-\sqrt{\beta}\sin\left(\frac{(y+1)k\pi}{K+1}\right)\right]\text{NB}(z;y,\mu)\biggr{|},$ (42)
where
$\alpha_{k}=\lambda\mu+\bar{\lambda}\bar{\mu}+2\sqrt{\lambda\mu\bar{\lambda}\bar{\mu}}\cos\left(\frac{k\pi}{K+1}\right)\text{ and }\beta=\frac{\lambda\bar{\mu}}{\mu\bar{\lambda}}.$
Proof: As Theorem 1 indicates, to analyze the predictability of this system with perfect observations, we require the derivation of 3 components: state probabilities for a lead time $L$, given an observed state $x$, stationary probabilities of states $\pi(\cdot)$, and the posterior distribution $\Pr(Z_{n}\mid X_{n}=x)$. The posterior distribution in this case is the sojourn time distribution conditioned on the queue size. For a Geo/Geo/1 queue with service probability $\mu$ and $x$ packets in the system, the sojourn time distribution is known as
$\Pr(Z_{n}\mid X_{n}=x)\sim\text{NB}(x+1,\mu),$ (43)
where NB denotes the negative binomial density function. As for the transient state probabilities $L$ timeslots ahead, they could be derived by raising the transition matrix to $L$ i.e. $P^{L}(x,\cdot)$ or using its another form derived in [36] given by:
$P^{L}(i,j)=\\ =\frac{1-\beta}{1-\beta^{K+1}}\beta^{j}+\frac{2}{K+1}\beta^{\frac{j-i}{2}}\sum_{k=1}^{K}\frac{\alpha_{k}^{L}}{\left[1-2\sqrt{\beta}\cos\left(\frac{k\pi}{K+1}\right)+\beta\right]}\\ \left[\sin\left(\frac{ik\pi}{K+1}\right)-\sqrt{\beta}\sin\left(\frac{(i+1)k\pi}{K+1}\right)\right]\\ \left[\sin\left(\frac{jk\pi}{K+1}\right)-\sqrt{\beta}\sin\left(\frac{(j+1)k\pi}{K+1}\right)\right].$ (44)
When the queue is stable, i.e. $\mu>\lambda$, the stationary probabilities are known as
$\pi(y)=\frac{1-\beta}{1-\beta^{K+1}}\beta^{y}.$ (45)
Then, we derive predictability by replacing Equations 43, 44, and 45 in Theorem 1 to complete the proof.
We propose an alternative approach to the aforementioned method for two reasons. Firstly, because of the possible computational demands, and secondly, because of the difficulty in determining the relationship between model parameters and predictability. Hence, we present this approximation in the next proposition.
Proposition 1
The predictability of the sojourn time of a Geo/Geo/1/K queue with a very large $K$ for $L$ time slots in future based on the observation of the queue size $x$ is approximated as
[…]
where
$\kappa=\frac{\sqrt{\lambda\mu\bar{\lambda}\bar{\mu}}}{\lambda\mu+\bar{\lambda}\bar{\mu}+2\sqrt{\lambda\mu\bar{\lambda}\bar{\mu}}}.$ (47)
Proof: see appendix B.
Compared to Theorem 3, using this approximation we can focus only on the terms that are critical in terms of influence on predictability. For instance, the term outside the integral which is raised to $L$. If it is reduced, as $L$ grows, predictability diminishes more rapidly. This implication necessitates understanding when the approximation is valid and incorporating that understanding into the problem at hand. For example, as indicated in the proof, the approximation assumes large value for $K$. We will evaluate this in more detail in the numeric section.
In the following section we focus on numerical evaluations of the predictability analysis methods in different scenarios and configurations.
V Numerical Evaluations
This section is dedicated to evaluating the proposed methods and examining the predictability in abstract systems in 3 subsections [^1] MATLAB code and reproducible results are available at https://github.com/samiemostafavi/predictability. First, we evaluate the spectral-based upperbound and approximations on predictability of the Geo/Geo/1/K queue. The second subsection is devoted to analyzing how dynamics of the queue affect its predictability and in the end multi-hop queues are analyzed when their observability is limited.
V-A Approximation and Upper Bound Validation
To assess the effectiveness of our proposed approximation method in Proposition 1 and the upper bound in Theorem 2, we analyze a Geo/Geo/1/K queue with default parameters $\mu=0.5$, $\rho=0.85$, and $K=128$. Figure 4 illustrates the posterior and marginal delay distributions given various observed queue lengths for this queue.
In Figure 5, we validate the approximation method by comparing it to the exact predictability derivation presented in Theorem 1. The main finding is that the accuracy of the approximation is significantly influenced by the disparity between the observed state’s forecast distribution and the marginal distribution. As illustrated in Figure 4, states that are closer to the marginal distribution display less accurate approximations in Figure 5, whereas states further from the marginal distribution (those above 16) demonstrate more accurate approximations.
Figure 4: Posterior and marginal distributions of a Geo/Geo/1/K queue with $\mu=0.5$, $\rho=0.85$, and $K=128$.
Figure 5: Predictability evaluation for different observed states of a Geo/Geo/1/K system with $\mu=0.5$, $\rho=0.85$, and $K=128$. The approximation method from Proposition 1 (dashed lines) is compared against the exact predictability derivation via Theorem 1 (solid lines).
Figure 6, shows the same setup as Figure 5, but with the full spectral-based upper bound in Theorem 2 (UB1). It is evident from the figure that unlike the approximation method, the upper bound is tighter for lower queue sizes. As the queue size grows, the distance between the exact curve and the bound increases. Due to poor performance, the spectral-gap based upper bound (UB2) is not involved in this analysis.
Figure 6: Predictability evaluation for different observed states of a Geo/Geo/1/K system with $\mu=0.5$, $\rho=0.85$, and $K=128$. The full spectral-based upper bound from Theorem 2 (dashed lines) is compared against the exact predictability derivation via Theorem 3 (solid lines).
In the next analysis, we apply state aggregation on the same queue. Figure 7 illustrates four levels of sequential and even state aggregations: no aggregation, 2-state aggregation, 4-state aggregation, and 8-state aggregation. The results show higher levels of aggregation lead to a less predictable system and more rapid decline of predictability.
Figure 7: Predictability of a Geo/Geo/1/K queue, under different state aggregation levels and $\rho=0.85$, $\mu=0.5$. Solid lines illustrate numerical derivation and dashed lines the approximation using Proposition 1.
The evaluation is illustrated as dashed lines in Figure 7 where the approximation closely follow the numerical derivation. However, it is important to note that the current approach involves an even aggregation of states. Future research could investigate uneven state aggregation to identify potential optimization strategies that might better balance the trade-off between observations’ overhead and predictability.
V-B Geo/Geo/1/K Dynamics and Predictability
To investigate the predictability in systems with Geo/Geo/1/K queues further, next, we examine the impact of the observed state and utilization factor ($\rho$). In the following analysis, we obtain predictability using Theorem 3 and Proposition 1.
Figure 8: Predictability of a single Geo/Geo/1 queue, under different configurations. Solid lines illustrate numerical derivation and dashed lines the approximation in Proposition 1.
Figure 8 demonstrates the predictability when $\mu=0.4$, $K=100$, and the observed state $x$ is either 3, 9, or 15 times greater than the expected state denoted by $\chi$. It is evident from the figure that the sojourn time is more predictable when there is a significant backlog or when the queue length is substantially different from the expected queue length. This implies that the queue remains predictable for a certain duration into the future, and we have sufficient information to predict the delay distribution for a period of time before it converges to the stationary delay distribution. In real systems, especially those operating in high utilization regimes where efficiency is critical, this analysis highlights the importance of monitoring the queue. High utilization often corresponds to more significant backlog situations, where predictability is maintained for longer durations.
In definition 2, we introduced the time frame during which the delay predictions remain distinct from the stationary distribution, referred to as epsilon-predictable horizon. Figure 9 illustrates the epsilon-predictable horizon of different Geo/Geo/1 queues. For queues with higher utilization, as predictability degrades more slowly, the forecasting system can sample the queue length less frequently and reduce the monitoring overhead. In a queue with higher arrival rate, but similar service rate, the stationary state distribution is wider and its expected queue length is larger, while the conditional distribution is unchanged. Therefore, the total variation difference between the forecast and marginal distributions decreases slower as we increase the lead time, resulting in a slower decline in predictability, as illustrated in the figures.
Figure 9: Epsilon-predictable horizon of Geo/Geo/1/K queuing system, under different configurations and $\epsilon=0.1$
Figure 9 illustrates epsilon-predictable horizon for Geo/Geo/1/K queues with $K=50$, highlighting its relationship with both the utilization factor and the service rate. In this figure, we observe that epsilon-predictable horizon increases in higher utilizations and lower service rates. Epsilon-predictable horizon is related to the state that results minimum distance between the forecast and marginal distributions, which is close to the expected queue length.
V-C Predictability of Multi-Hop Geo/Geo/1 Queues
In this part, we evaluate 2 multi-hop scenarios under different levels of observability. In the first case, illustrated in Figure 10, we evaluate predictability for a system that consists of 5 tandem Geo/Geo/1/K queues, all with the same service rate $\mu=0.4$ and highly congested with the observed state of $15\chi$. We consider 3 different configurations: observations that include all 5 queues, observations including only the first 3 queues, and observations including only the first queue. We used Equations 37 and 38to derive exact predictability curves shown via solid lines. As was expected, predictability decreases as our ability to observe the system declines. Then the dashed line curves show the upper bounds derived using Lemma 5 and Theorem 3 (UB). In the first case where all queues are observable, we derived the bound by summing up the predictability of all hops. However, when it comes to the case with only one observable queue, the bound equals the predictability of that queue. This is because if a component is not observable, its predictability is zero and will not contribute to the sum in Equation 39. Therefore, we have $D_{n}(L)\leq D^{(1)}_{n}(L)$, which constitutes the upper bound for the partially observable system.
Figure 10: Predictability of a 5-hop tandem Geo/Geo/1 queuing system under different observability levels. All queues have the same service rate of $\mu=0.4$ and $\lambda=0.32$
The next scenario we analyze in multi-hop networks involves a queue that becomes a bottleneck and how its observability affects predictability. Since this queue will contribute more to the end-to-end delay, if it is not observable, we expect the predictability to decrease significantly. In a setup, we examine 3 tandem Geo/Geo/1/K queues with an arrival rate of $\lambda=0.34$, where the middle queue has a lower service rate, i.e., the service rates are $\mu^{(1)}=0.4$, $\mu^{(2)}=0.38$, and $\mu^{(3)}=0.4$ respectively. Suppose $\chi^{(m)}$ as the expected queue length of hop $m$. All three queues are again highly congested with the observed states of 10 times their expected queue size. As shown in Figure 11, we consider 4 different configurations: the observations include all 3 queues, the observations include only the first queue, the observations include only the last queue, and the observation include only the bottleneck queue. Note that in tandem M/M/1 queues, when the system is in equilibrium, regardless of the service rates, all queues will have the same arrival rate as the first queue [37]. Using that, Equation 37, and Equation 38, we derive exact predictability curves shown via solid lines. The figure indicates that the observation of one of the less utilized queues provides similar benefits in terms of predictability. However, observing these queues alone results in significantly lower predictability compared to the scenario in which only the bottleneck queue is observable. The predictability of the latter case is very close to that of complete observability. Similar to Figure 10, the dashed lines represent the upper bound derived using Lemma 5 and Theorem 3 (UB) and it is less tight when the predictability of all hops is summed up. Note that $\chi^{(1)}=\chi^{(3)}$ due to similar arrival and service rates.
Figure 11: Predictability of a 3-hop tandem Geo/Geo/1/K queuing system with a bottleneck queue in the middle. Service rates are $0.4$, $0.38$, and $0.4$ respectively. Arrival rate: $0.34$.
VI Conclusions
In this paper, we presented a definition of performance predictability for systems with Markovian conditions, with partial and imperfect observability. We applied this framework to Geo/Geo/1 queuing models in both single-hop and multi-hop contexts. We illustrated the practical implications of studying predictability in these systems. Our evaluations highlight the effectiveness of the proposed definitions in modeling crucial trade-offs in future communication systems incorporating QoS predictors. For example, we illustrated the overhead-predictability tradeoff and demonstrated how a lack of observability can deteriorate predictability.
In multi-hop networking, we showed both exact and upper bounds on the predictability, with reduced monitoring resources compared to observing all queues, showed monitoring the bottleneck queue is crucial for maintaining predictability. Additionally, state aggregation is a viable option to manage the overhead of observations in large-scale networks.
The behavior of the underlying Markov processes greatly impacts the predictability of the system. Drawing on the literature on Markov chain mixing times, including spectral gap analysis, provides insights into how random system conditions transition. We examined the connection between system predictability and Markov chain behavior by employing an upper limit on predictability derived from spectral analysis methods. Future research can investigate conductance analysis to further improve our comprehension of Markov chain dynamics. A possible direction for future research is to identify or detect hidden Markov models in the environment with which the communication network interacts. Incorporating these models into predictability analysis will offer insights into the feasibility and timescales of QoS predictions.
In future communication systems where performance prediction is essential, it is important to improve predictability efficiently. This involves enhancing observability and directing monitoring efforts towards critical points within the network, such as performance bottleneck nodes or queues. Additionally, it is necessary to identify whether the environment or application’s behavior is predictable and subsequently learn it for performance prediction. By incorporating these insights, more efficient communication systems can be achieved.
Appendix A Proof of Theorem 2
In this proof, our approach is to bound total variation distance using chi-squared distance. For arbitrary probability distributions $p$ and $q$, according to Pinsker’s inequality, total variation distance is upper bounded as
$\lVert p-q\rVert_{\text{TV}}\leq\sqrt{\frac{1}{2}\chi^{2}(p,q)}$ (A.1)
where $\chi^{2}(p,q)$ denotes chi-squared distance and is obtained
$\chi^{2}(p,q)=\sum_{z\in\mathcal{Z}}\frac{(p(z)-q(z))^{2}}{q(z)}.$ (A.2)
Therefore, we begin by bounding chi-squared distance between forecast and marginal distributions which can be derived as
$\chi^{2}(\Pr(Z_{n+L}\mid X_{n}=x),\Pr(Z_{n+L}))=\\ \sum_{z\in\mathcal{Z}}\frac{\left(\sum_{y\in\mathcal{X}}\left(P^{L}(x,y)-\pi(y)\right)r_{y}(z)\right)^{2}}{Pr(Z_{n+L})}.$ (A.3)
We use the following equations in our proof from the spectral theorem in Markov chain mixing time [38]:
$\frac{P^{L}(x,y)}{\pi(y)}=1+\sum_{j=2}^{K}f_{j}(x)f_{j}(y)\lambda_{j}^{L},$ (A.4)
$\sum_{i=2}^{n}f_{i}(x)f_{i}\left(x^{\prime}\right)=\frac{\mathbf{1}\left(x=x^{\prime}\right)}{\pi(x)}-1.$ (A.5)
First, we substitute $P^{L}(x,y)-\pi(y)$ from above in A.3 and obtain
$\chi^{2}(\Pr(Z_{n+L}\mid X_{n}=x),\Pr(Z_{n+L}))=\\ \sum_{z}\frac{\left(\sum_{y}\sum_{j=2}^{|\mathcal{X}|}f_{j}(x)f_{j}(y)\lambda_{j}^{L}\pi(y)r_{y}(z)\right)^{2}}{\sum_{y}{\pi(y)r_{y}(z)}}.$ (A.6)
Then we apply Cauchy-Schwarz inequality as follows
$\left(\sum_{y}\sum_{j=2}^{|\mathcal{X}|}f_{j}(x)f_{j}(y)\lambda_{j}^{L}\pi(y)r_{y}(z)\right)^{2}=\\ \left(\sum_{j=2}^{|\mathcal{X}|}f_{j}(x)\lambda_{j}^{L}\sum_{y}f_{j}(y)\pi(y)r_{y}(z)\right)^{2}\leq\left(\sum_{j=2}^{|\mathcal{X}|}f^{2}_{j}(x)\lambda^{2L}_{j}\right)\cdot\\ \left(\sum_{j=2}^{|\mathcal{X}|}\sum_{y,y^{\prime}}f_{j}(y)\pi(y)r_{y}(z)f_{j}(y^{\prime})\pi(y^{\prime})r_{y^{\prime}}(z)\right)\\ \leq\lambda^{2L}_{*}\left(\sum_{j=2}^{|\mathcal{X}|}f^{2}_{j}(x)\right)\cdot\\ \left(\sum_{y,y^{\prime}}\pi(y)r_{y}(z)\pi(y^{\prime})r_{y^{\prime}}(z)\sum_{j=2}^{|\mathcal{X}|}f_{j}(y)f_{j}(y^{\prime})\right)\\ =\lambda^{2L}_{*}\left(\frac{1}{\pi(x)}-1\right)\cdot\\ \left(\sum_{y,y^{\prime}}\pi(y)r_{y}(z)\pi(y^{\prime})r_{y^{\prime}}(z)\left(\frac{\mathbf{1}(y=y^{\prime})}{\pi(y)}-1\right)\right),$ (A.7)
where $\mathbf{1}()$ denotes the unit function.
$\chi^{2}(\Pr(Z_{n+L}\mid X_{n}=x),\Pr(Z_{n+L}))\leq\lambda^{2L}_{*}\left(\frac{1}{\pi(x)}-1\right)\cdot\\ \sum_{z}\frac{\sum_{y,y^{\prime}}\pi(y)r_{y}(z)\pi(y^{\prime})r_{y^{\prime}}(z)\left(\frac{\mathbf{1}(y=y^{\prime})}{\pi(y)}-1\right)}{\sum_{y}{\pi(y)r_{y}(z)}}$ (A.8)
and for the right term we simplify
$\sum_{z}\frac{\sum_{y,y^{\prime}}\pi(y)r_{y}(z)\pi(y^{\prime})r_{y^{\prime}}(z)\left(\frac{\mathbf{1}(y=y^{\prime})}{\pi(y)}-1\right)}{\sum_{y}{\pi(y)r_{y}(z)}}=\\ \sum_{z}\frac{\sum_{y}\pi(y)r^{2}_{y}(z)}{\sum_{y}{\pi(y)r_{y}(z)}}-1$ (A.9)
$\chi^{2}(\Pr(Z_{n+L}\mid X_{n}=x),\Pr(Z_{n+L}))\leq\\ \lambda^{2L}_{*}\left(\frac{1}{\pi(x)}-1\right)\left(\sum_{z}\frac{\sum_{y}\pi(y)r^{2}_{y}(z)}{\sum_{y}{\pi(y)r_{y}(z)}}-1\right),$ (A.10)
Incorporate this into A.1 we get
$D_{n}(L)=\lVert\Pr(Z_{n+L}\mid X_{n}=x)-\Pr(Z_{n+L}))\rVert_{\text{TV}}\leq\\ \frac{1}{\sqrt{2}}\lambda^{L}_{*}\left(\frac{1}{\pi(x)}-1\right)^{1/2}\left(\sum_{z}\frac{\sum_{y}\pi(y)r^{2}_{y}(z)}{\sum_{y}{\pi(y)r_{y}(z)}}-1\right)^{1/2}.$ (A.11)
Appendix B Proof of Proposition 1
We start from Theorem 3, Equation 42 where in the inner sum over $k$, the term $\alpha^{L}_{k}$ converges to zero, while other terms remain nonzero, but only when $L$ is large, and except when $k/(K+1)$ is small. Therefore, we simplify the rest of the terms for the case when $k/(K+1)$ is small. As for the $\sin$ terms, as $k/(K+1)$ is small, we approximate
$\sin\left(\frac{(x+1)k\pi}{K+1}\right)\approx\sin\left(\frac{xk\pi}{K+1}\right),\\ \sin\left(\frac{(y+1)k\pi}{K+1}\right)\approx\sin\left(\frac{yk\pi}{K+1}\right)$ (B.1)
We can simplify Theorem 3 as
$D_{n}(L)=\frac{1}{2}\sum_{z}\biggl{|}\sum_{y=1}^{K}\frac{2}{K+1}\beta^{\frac{y-x}{2}}(1-\sqrt{\beta})^{2}\\ \sum_{k=1}^{K}\frac{\left(\lambda\mu+\bar{\lambda}\bar{\mu}+2\sqrt{\lambda\mu\bar{\lambda}\bar{\mu}}\cos(\frac{k\pi}{K+1})\right)^{L}}{\left[1-2\sqrt{\beta}\cos\left(\frac{k\pi}{K+1}\right)+\beta\right]}\\ \sin\left(\frac{xk\pi}{K+1}\right)\sin\left(\frac{yk\pi}{K+1}\right)\text{NB}(z;y,\mu)\biggr{|}.$ (B.2)
We continue the approximation with the summation over $y$:
$\sum_{y=1}^{K}\beta^{y/2}\sin{\left(\frac{yk\pi}{K+1}\right)}\text{NB}(z;y,\mu)\approx\\ \sqrt{\beta}\frac{\sin{\left(\frac{k\pi}{K+1}\right)}}{1-2\sqrt{\beta}\cos{\left(\frac{k\pi}{K+1}\right)}+\beta}\text{NB}(z;1,\mu)$ (B.3)
Since $\beta^{y/2}$ decays exponentially, the dominant contributions come from the lower values of $y$. Therefore, we substituted the NB term with $\text{NB}(z;1,\mu)$. The remaining terms inside the summation are easy to show that for large $K$, and $\beta<1$, can be approximated accordingly.
Next we simplyfy $\alpha_{k}^{L}$ term using the binomial approximation for large $L$:
$(a+b\cos(c))^{L}\approx(a+b)^{L}\left(1-\frac{bc^{2}}{2(a+b)}\right)^{L}\approx\\ (a+b)^{L}e^{-L\frac{bc^{2}}{2(a+b)}}$
Substitute $\alpha_{k}^{L}$ with above approximation will give us:
$\left(\lambda\mu+\bar{\lambda}\bar{\mu}+2\sqrt{\lambda\mu\bar{\lambda}\bar{\mu}}\cos(\frac{k\pi}{K+1})\right)^{L}\approx\\ \left(\lambda\mu+\bar{\lambda}\bar{\mu}+2\sqrt{\lambda\mu\bar{\lambda}\bar{\mu}}\right)^{L}e^{-L\kappa(k\pi/(K+1))^{2}}$ (B.4)
where
$\kappa=\frac{\sqrt{\lambda\mu\bar{\lambda}\bar{\mu}}}{\lambda\mu+\bar{\lambda}\bar{\mu}+2\sqrt{\lambda\mu\bar{\lambda}\bar{\mu}}}$
Next, since $K$ is large, we convert the sum $\sum_{k}$ to integral by substituting $k\pi/(K+1)$ with $r$ as
[…]
Since the integral and the rest of the terms are positive, we remove the absolute value operator and move $\sum_{z}$ inside to the only term dependent on $z$, which is the NB term and since it is a pmf, the summation results $1$.
Finally, by combining all approximations we reach the final term as
[…]